1. Combining chlorine gas and oxygen gas can produce dichlorine heptoxide.  If 87.8 L of oxygen is combined with 62.5 L of chlorine, how many grams of dichlorine heptoxide will be produced?
1
Expert's answer
2021-05-25T07:50:13-0400

2Cl2(g)+7O2(g)"\to" 2Cl2O7(g)

Moles of Cl2=(87.8*1)/24=3.658moles

Moles of O2=(62.5*1)/24=2.604moles

If 7=2.604moles

"\therefore" 2=(2*2.604)/7=0.74405moles

RMM of Cl2O7=182.901g/mol

Mass of Cl2O7=(182.901*0.74405)

=136.09g